OpenStudy (alexh107):

Difference quotient for f(x) = 4x

Directrix (directrix):

f(x + h) - f(x) ------------ = ? h

OpenStudy (alexh107):

I have the formula in front of me I just don't know how to apply it

Directrix (directrix):

In f(x) = 4x, the function f takes x and multiplies it by 4. In f(x+h), the function f will take (x + h) and replace x by that to get f(x + h) = 4*(x + h) Agree?

OpenStudy (alexh107):

I think so. What's confusing me is the h.

Directrix (directrix):

f(x + h) - f(x) 4x + 4h - 4x 4h ------------ = -------------- = ---- = 4 h h h

Directrix (directrix):

Don't think too much about the h. Just ask yourself, what is f doing to what is within the ( ) of f(x). In this case f is taking x and multiplying it by 4. So, when (x + h) gets with f as in f(x+h), the the (x+h) is multiplied by 4.

OpenStudy (alexh107):

Oh okay. I think I get it better now. Thank you for your help!

Directrix (directrix):

When you get on down the line in this class, the h will take on more significance. But if f(x) = 4*x, f(x+h) will be 4*(x+h). Just think of x in f(x) as a placeholder of sorts.

OpenStudy (alexh107):

Yea that makes sense

Directrix (directrix):

The first time I studied this, I was confused.

OpenStudy (alexh107):

Yeah, I think it just takes some practice to get it down completely

OpenStudy (alexh107):

But thank you for explaining it!

Directrix (directrix):

You are welcome.